OpenStudy (anonymous):

The inner and the outer radii of a cylindrical pipe are 5cm and 4cm respectively.Find the area of cross section of the pipe(Take pi=3.14).

ganeshie8 (ganeshie8):

inner radius = 5 outer radius = 4 how can inner radius be greater than outer radius ? question looks wrong lol

OpenStudy (anonymous):

is it inner 4 ???

OpenStudy (anonymous):

I THINK ITS PRINTED WRONG! inner=4 and outr=5

OpenStudy (anonymous):

plz. help anyone!

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

it just the area of inside circle. are of circle = \(\pi r^2 = 3.14 * 4^2\)

OpenStudy (anonymous):

pi(Routersquared- Routersquared)

OpenStudy (anonymous):

i got it! pi(R sq. - R sq. ) = pi (R+r)* (R-r) = 3.14* (5+4)* (5-4) = 3.14*20*1 = 62.80

ganeshie8 (ganeshie8):

hmm how u got that ? cross section of pipe is just the inner circle area, right ? it should be pi(r^2) = 3.14(4)(4) = 50.24